Opened 10 years ago

Closed 9 years ago

Last modified 3 years ago

#255 closed flight processing (fixed)

GB08/06, flight day 194/2009, Bleaklow

Reported by: knpa Owned by: knpa
Priority: alpha 4 low Milestone: 2009 Data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_06-2009_194_Bleaklow

Data arrived from ARSF via SATA disk on 17/7/09

Scientific objective: Peatland monitoring

Priority: a4l

PI: Julia McMorrow

Sensors:

  • Eagle (13/11/09)
  • Hawk (13/11/09)
  • LIDAR (22/06/10)
  • Digital Photography (10/02/2010)

Change History (35)

comment:1 Changed 10 years ago by knpa

  • Description modified (diff)

comment:2 Changed 10 years ago by knpa

  • Owner mggr deleted

comment:3 Changed 10 years ago by mggr

  • Component changed from ARSF to Processing: general

comment:4 Changed 10 years ago by mggr

  • Milestone set to 2009 Data processing completion

comment:5 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:6 Changed 10 years ago by mggr

  • Description modified (diff)

Has lidar data too.

comment:7 Changed 10 years ago by knpa

Applanix processing completed

comment:8 Changed 10 years ago by knpa

E+H Processing Completed

comment:9 Changed 10 years ago by knpa

Lev1 files (and overview screen-shot) made available to PI via ftp.

comment:10 Changed 10 years ago by knpa

non-correct lev1 images deleted and correct ones renamed (but may need further renaming)

comment:11 Changed 10 years ago by knpa

e-mailed PI to ask if they would like a pre-vector delivery

comment:12 Changed 10 years ago by mark1

Vectors have arrived.

comment:13 Changed 10 years ago by emca

  • Owner changed from knpa to emca
  • Status changed from assigned to new

comment:14 Changed 10 years ago by emca

Delivery check complete
Data looks good, made some changes to read_me and lev1 file names

comment:15 Changed 10 years ago by emca

  • Owner emca deleted

Eagle and Hawk data sent to J McMorrow 13/11/09

comment:16 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:17 Changed 10 years ago by mark1

  • Description modified (diff)

Photography has been dispatched to PI on 10th February 2010

comment:18 Changed 10 years ago by mark1

  • Description modified (diff)

comment:19 Changed 10 years ago by knpa

  • Description modified (diff)

Started LIDAR processing.

comment:20 Changed 10 years ago by knpa

ipas appears to have been done already.

comment:21 Changed 10 years ago by knpa

Large holes in lidar data for most flightlines

comment:22 Changed 10 years ago by knpa

incorrect directions on logsheet

comment:23 Changed 10 years ago by knpa

Flightlines processed, currently QCing.

Vertical discrepancies between flightlines seem acceptable, despite 2 or 3 large ones (over 10cm).

Can see no evidence of roll for this project.

comment:24 Changed 10 years ago by knpa

lines line up horizontally also.

comment:25 Changed 9 years ago by mark1

Lidar delivery has been checked and looks OK. Ready to deliver.

comment:26 Changed 9 years ago by knpa

NOTE This project was reprocessed due to incorrect min-range settings cutting off some of the data so at the moment there are two lots of flightlines in the workspace.

comment:27 Changed 9 years ago by knpa

  • Description modified (diff)

LIDAR delivered today (22/06/10) to Julia McMorrow at Manchester.

comment:28 Changed 9 years ago by knpa

LIDAR still needs to be rsynced (no room currently)

comment:29 Changed 9 years ago by knpa

Have rsynced data for all sensors.

comment:30 Changed 9 years ago by knpa

Beginning to archive.

comment:31 Changed 9 years ago by knpa

made available to NEODC as tarball

comment:32 Changed 9 years ago by knpa

Downloaded and backed-up by NEODC.

Moved to archive disk.

comment:33 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleting workspace version and closing ticket.

comment:34 Changed 3 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:35 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.